OpenStudy (kirbykirby):

If the sequence|{An}| converges, does {An} converge?If the sequence|{An}| converges, does {An} converge?@Mathematics

OpenStudy (kirbykirby):

(the first An is in absolute value)

OpenStudy (agreene):

So, are you asking the inverse of the test for absolute convergence... or something?

OpenStudy (agreene):

\[\sum A_n\] is absolutely convergent iff \[\sum |A_n| < \infty\]

OpenStudy (kirbykirby):

no not series, the sequence An